Two runners approaching each other on a straight track have constant speeds of 4.50 m/s and 3.50 m/s, respectively, when they are 100 m apart. How long will it take for the runners to meet, and at what position will they meet if they maintain these speeds?
Answer by TimothyLamb(4379)

speed of runners relative to each other:
4.50 + 3.50 = 8 m/s
---
s = d/t
t = d/s
t = 100/8
t = 12.5 seconds
---
position of fast runner at the meet:
d = st
d = 4.5 * 12.5
d = 56.25 m from his initial position
---
position of slow runner at the meet:
d = st
d = 3.5 * 12.5
d = 43.75 m from his initial position
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---